xy^2 +5 = 9x−3y^2


1
Expert's answer
2021-05-07T10:32:02-0400

xy^2 +5 = 9x−3y^2

Differentiation with respect to x

"x\times2y\frac {dy}{dx}+6y\frac{dy}{dx}=9-y^2"

"(2xy+6y)\frac{dy}{dx}=9-y^2"

"\frac{dy}{dx}=\frac{9-y^2}{2xy+6y}"